# Least multivariate Chebyshev polynomials on diagonally-determined domains

### v2

@@ 4. The dual framework and signatures
In this section we consider the dual problem of (1.3). Our goal is to show that, in the special case of diagonally-determined domains, the dual problem has a closed-form solution. Moreover, we will construct a dual solution that is atomic (discrete), and supported on $n+1$ points.
@@ 4.1. The dual problem
With $\Omega$ being a domain in ${\mathbb{R}}^{d}$, let $C(\Omega)$ be the space of continuous functions on $\Omega$, and let $C(\Omega)^{*}$ be the dual space of $C(\Omega)$. For compact $\Omega$, one has the following, strong duality result for problem (1.3).
@@
$\|P^{*}\|_{\Omega}=\min_{Q\in\Pi_{n-1}^{d}}\left\|\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}+Q({\mathbf{x}})\right\|_{\Omega},$
@@
Recall the strong duality result for the classical Chebyshev approximation problem as given in the paper by Rivlin and Shapiro [13, Corollary 2]. [^1] In [13, Corollary 2], the objective is in fact given as $\sup|L(f)|$, but the absolute value may be omitted w.l.o.g., since $L$ is feasible for the dual problem if and only if $-L$ is feasible.
@@
For $f({\mathbf{x}})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}$, any optimal solution of problem (4.1) is also optimal for the dual problem in (4.3).
@@
Let $L^{*}$ denote an optimal solution of problem (4.1). Then $L^{*}$ is feasible for the dual (maximization) problem in (4.3). Moreover,
$L^{*}(f)=L^{*}\left(\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}\right)=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}L^{*}({\mathbf{x}}^{{\alpha}})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}\gamma^{*}=\gamma^{*}=\|P^{*}\|_{\Omega}.$
This yields the result. ∎
@@ 4.2. Atomic dual solutions and signatures
The first statement follows immediately from Theorem 4.4, applied to an optimal solution of problem (4.1), and using $V=\Pi_{n}^{d}$. It remains to show that the points $\omega_{1},\ldots,\omega_{r}$ are extremal points of any optimal solution $P^{*}$ to problem (1.3). Now let $L^{*}$ denote an optimal solution of problem (4.1), so that
@@
Since $\sum_{i=1}^{r}|\tau_{i}|=1$, $\|P^{*}\|_{\Omega}$ is a weighted average of the values $P^{*}(\omega_{i})$ with $\tau_{i}>0$ and $-P^{*}(\omega_{i})$ with $\tau_{i}<0$. As a consequence, we have
@@
A signature with finite support $S\subset\Omega$ is simply a (partition) function from $S$ to {±$\{\pm 1\}$. A signature $\sigma$ with support $S$ is said to be extremal for a subspace $V\subset C(\Omega)$ if there exist weights $\lambda_{\omega}>0$, $\omega\in S$, such that $\sum_{\omega\in S}\lambda_{\omega}\sigma(\omega)v(\omega)=0$ for all $v\in V$. A signature $\sigma$ with support $S$ is said to be associated with a function $g\in C(\Omega)$ if $S$ is included in the set $\{\omega\in\Omega:|g(\omega)|=\|g\|_{\Omega}\}$ of extremal points of $g$ and if $\sigma(\omega)=\operatorname{sgn}(g(\omega))$ for all $\omega\in S$.
Thus, the result of Corollary 4.5 may be restated as the existence of a signature with support $S=\{\omega_{1},\ldots,\omega_{r}\}$ where $r\leq\mbox{dim}(\Pi_{n}^{d})$. This signature is extremal for $\Pi_{n-1}^{d}$, and associated with every optimal solution of problem (1.3). Therefore, we will simply refer to an optimal signature for problem (1.3). Formally we have the following result.
@@
4.3. Signatures for diagonally-determined domains
@@
$P^{*}({\mathbf{x}})=\left(\frac{b-a}{2}\right)^{n}\frac{1}{2^{n-1}}T_{n}\left(-1+2\,\frac{\langle{\mathbf{v}},{\mathbf{x}}\rangle-a}{b-a}\right).$
@@
$-1+2\,\frac{\langle{\mathbf{v}},\omega_{j}\rangle-a}{b-a}=\xi_{j}\text{ for }j=1,\ldots,n+1.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{*}=\frac{1}{n}\sum_{i=1}^{n+1}(-1)^{i+1}\lambda_{i}L_{\omega_{i}}.}$
We claim that $L^{*}$ is an optimal solution of the dual problem (4.1). We first verify feasibility. We will show that $L^{*}$ vanishes on $\Pi^{d}_{n-1}$, as required, by using the Gauss-Lobatto-Chebyshev quadrature formula, which implies, for $k<n$:
[…]
@@
$=$ $\frac{1}{n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}\frac{1}{2}[T_{|k_{1}-k_{2}|}(\xi_{j})+T_{k_{1}+k_{2}}(\xi_{j})]$
@@ 5. Detecting the diagonally-determined property
The quadratic module of $g_{1},\ldots,g_{m}$, denoted by $\mathcal{Q}(g_{1},\ldots,g_{m})$, is defined as all polynomials of the form
@@
We assume that $\mathcal{Q}(g_{1},\ldots,g_{m})$ satisfies the Archimedean condition, meaning that there exists a $p\in\mathcal{Q}(g_{1},\ldots,g_{m})$ such that
$\left\{{\mathbf{x}}\in\mathbb{R}^{d}\;:\;p({\mathbf{x}})\geq 0\right\}\mbox{ is compact.}$.
@@
We also introduce the concept of the truncated quadratic module of degree $k\in\mathbb{N}_{0}$, denoted by $\mathcal{Q}(g_{1},\ldots,g_{m})_{k}$, as the subset of $\mathcal{Q}(g_{1},\ldots,g_{m})$ where each term on the right-hand-side of (5.2) has degree at most $k$, i.e. deg$(\sigma_{0})\leq k$, and deg$(\sigma_{j}g_{j})\leq k$ for all $j=1,\ldots,m$. The membership problem for $\mathcal{Q}(g_{1},\ldots,g_{m})_{k}$ may be reformulated as a semidefinite programming feasibility problem; see Lasserre [6], and the references therein.
@@
We assume that we now know $[a,b]=\mathrm{diag}(\Omega)$, and that $a<0<b$. The latter assumption is without loss of generality, due to Lemma 2.7. Then, to detect whether the second property holds in Definition 2.1, consider the optimization problem:
@@
[…]
@@
Assume $\Omega$ is of the form (5.1), that Assumption 5.1 holds, and that $\mathrm{diag}(\Omega)=[a,b]$ where $a<0<b$. Then one has $\rho_{k}\in[0,1]$ for all $k$. Moreover, if $\Omega$ is diagonally-determined, then the sequence $\{\rho_{k}\}$ converges monotonically to zero. If $\rho_{k}=0$ for some finite value of $k$, then the optimal ${\mathbf{v}}^{\prime}\in\mathbb{R}^{d}$ of (5.4) certifies that $\Omega$ is diagonally-determined with vector ${\mathbf{v}}={\mathbf{v}}^{\prime}$ and diagonal $[a,b]$.
@@
where we again use that $a<0<b$. Thus, ${\mathbf{x}}\mapsto\langle(1-\varepsilon){\mathbf{v}},{\mathbf{x}}\rangle-a\in\mathcal{Q}(g_{1},\ldots,g_{m})$ and ${\mathbf{x}}\mapsto b-\langle(1-\varepsilon){\mathbf{v}},{\mathbf{x}}\rangle\in\mathcal{Q}(g_{1},\ldots,g_{m})$, by Theorem 5.2. This implies that, for sufficiently large $k^{*}\in\mathbb{N}_{0}$, ${\mathbf{v}}^{\prime}=(1-\varepsilon){\mathbf{v}}$ is feasible for (5.4) with $k=k^{*}$. Thus $\rho_{k^{*}}\leq 1-\langle(1-\varepsilon){\mathbf{v}},\mathbf{1}\rangle=\varepsilon$, and therefore the sequence $\{\rho_{k}\}$ converges monotonically to zero. ∎
@@
By Proposition 2.3, and its proof, we know that $\Omega$ is diagonally-determined, with diagonal $[a,b]=[-1,1]$ and vector ${\mathbf{v}}=\frac{1}{d}\mathbf{1}$. We may describe $\Omega$ as a semi-algebraic set in the form (5.1) by using the polynomials $g_{j}({\mathbf{x}})=1-x_{j}^{2}$ for $j=1,\ldots,d$, i.e. $m=d$. The next step is to find the diagonal of $\Omega$ by finding the roots of the univariate polynomials $p_{j}(t):=g_{j}(t\mathbf{1})=1-t^{2}$ for $j=1,\ldots,d$. These polynomials are the same for all $j$ and nonnegative on the interval $[-1,1]$, so that we obtain $[a,b]=[-1,1]$. The next step is to solve (5.4) for increasing values of $k\in\mathbb{N}$. Using the identity
$1\pm x=\frac{1}{2}\left((1\pm x)^{2}+1-x^{2}\right),\quad(x\in\mathbb{R}),$
@@
$\langle{\mathbf{v}},{\mathbf{x}}\rangle-a=\frac{1}{2d}\sum_{j=1}^{d}\left((1+x_{j})^{2}+g_{j}({\mathbf{x}})\right).$
@@ 6. Conclusion and discussion
[…]
@@
A second point to mention, is that we have not explored the geometric and invariance properties of diagonally-determined domains to any great extent in this paper, and this deserves to be investigated further. The same may be said for the decision problem of whether a given domain is diagonally-determined. If $\Omega$ is a compact semi-algebraic set, one may formulate an algorithmic procedure to decide this question, as we did in Section 5. An interesting special case is the (complexity) question: may one decide in polynomial time is a given polyhedron is diagonally-determined?

### v3

@@ 4. The dual framework and signatures
In this section we consider the dual problem of (1.3). Our goal is to show that, in the special case of diagonally-determined sets, the dual problem has a closed-form solution. Moreover, we will construct a dual solution that is atomic (discrete), and supported on $n+1$ points.
@@ 4.1. The dual problem
With $\Omega$ being a set in ${\mathbb{R}}^{d}$, let $C(\Omega)$ be the space of continuous functions on $\Omega$, and let $C(\Omega)^{*}$ be the dual space of $C(\Omega)$. For compact $\Omega$, one has the following, strong duality result for problem (1.3).
@@
$\|P^{*}\|_{\Omega}=\min_{Q\in\Pi_{n-1}^{d}}\left\|\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}+Q({\mathbf{x}})\right\|_{\Omega}=\min_{Q\in\Pi_{n-1}^{d}}\left\|\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}-Q({\mathbf{x}})\right\|_{\Omega},$
@@
Recall the strong duality result for the classical Chebyshev approximation problem as given in the paper by Rivlin and Shapiro [13, Corollary 2]. [^1] In [13, Corollary 2], the objective is in fact given as $\sup|L(f)|$, but the absolute value may be omitted without loss of generality, since $L$ is feasible for the dual problem if and only if L-L is feasible.
@@
For $f({\mathbf{x}})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}$ with $a^{*}_{\alpha}$ as defined in (4.2), and $\Omega$ compact, any optimal solution of problem (4.1) is also optimal for the dual problem in (4.3).
@@
Let $L^{*}$ denote an optimal solution of problem (4.1). Then $L^{*}$ is feasible for the dual (maximization) problem in (4.3). Using that $\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{{\alpha}}=1$, one has
$L^{*}(f)=L^{*}\left(\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}\right)=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}L^{*}({\mathbf{x}}^{{\alpha}})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}\gamma^{*}=\gamma^{*}=\|P^{*}\|_{\Omega},$
where the last equality is the strong duality relation in Theorem 4.1. Since the optimal value of the primal problem (4.3) is also $\|P^{*}\|_{\Omega}$ when $f({\mathbf{x}})=\sum_{{\alpha}\in{\mathbb{N}}^{d}_{n}}a^{*}_{\alpha}{\mathbf{x}}^{{\alpha}}$, it follows that $L^{*}$ is an optimal solution for the dual problem in (4.3), as required. ∎
@@ 4.2. Atomic dual solutions and signatures
Let $L^{*}$ denote an optimal solution of problem (4.1). The first statement in the corollary now follows immediately from Theorem 4.4, applied to $L^{*}$, and using $V=\Pi_{n}^{d}$, so that
@@
It remains to show that the points $\omega_{1},\ldots,\omega_{r}$ are extremal points of any optimal solution $P^{*}$ to problem (1.3).
@@
Since $\sum_{i=1}^{r}|\tau_{i}|=1$, $\|P^{*}\|_{\Omega}$ is a weighted average of the values $P^{*}(\omega_{i})$ with $\tau_{i}>0$ and $P^{*}(\omega_{i})$ with $\tau_{i}<0$. As a consequence, we have
@@
A signature with finite support $S\subset\Omega$ is simply a (partition) function from $S$ to $\{\pm 1\}$. A signature $\sigma$ with support $S$ is said to be extremal for a subspace $V\subset C(\Omega)$ if there exist weights $\lambda_{\omega}>0$, $\omega\in S$, such that $\sum_{\omega\in S}\lambda_{\omega}\sigma(\omega)v(\omega)=0$ for all $v\in V$. A signature $\sigma$ with support $S$ is said to be associated with a function $g\in C(\Omega)$ if $S$ is included in the set $\{\omega\in\Omega:|g(\omega)|=\|g\|_{\Omega}\}$ of extremal points of $g$ and if $\sigma(\omega)=\operatorname{sgn}(g(\omega))$ for all $\omega\in S$.
Thus, the result of Corollary 4.5 may be restated as the existence of a signature with support $S=\{\omega_{1},\ldots,\omega_{r}\}$ where $r\leq\mbox{dim}(\Pi_{n}^{d})$, so that this signature is extremal for $\Pi_{n-1}^{d}$, and associated with every optimal solution of problem (1.3). Therefore, we will simply refer to an optimal signature for problem (1.3). Formally we have the following result.
@@
4.3. Signatures for diagonally-determined sets
@@
${\mathbf{x}}\mapsto P^{*}({\mathbf{x}})=\left(\frac{b-a}{2}\right)^{n}\frac{1}{2^{n-1}}T_{n}\left(-1+2\,\frac{\langle{\mathbf{v}},{\mathbf{x}}\rangle-a}{b-a}\right).$
@@
$1+2\,\frac{\langle{\mathbf{v}},\omega_{j}\rangle-a}{b-a}=\xi_{j}\text{ for }j=1,\ldots,n+1.$
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}L^{*}=\frac{1}{n}\sum_{i=1}^{n+1}(-1)^{i+1}\lambda_{i}L_{\omega_{i}},$}
so that $\|L^{*}\|_{C(\Omega)^{*}}=\frac{1}{n}\sum_{i=1}^{n+1}|(-1)^{i+1}\lambda_{i}|=1$, by construction.
We claim that $L^{*}$ is an optimal solution of the dual problem (4.1). We first verify feasibility. We will show that $L^{*}$ vanishes on $\Pi^{d}_{n-1}$, as required, by using the Gauss-Lobatto-Chebyshev quadrature formula (see e.g. [12, Exercise 1.5.29]), which implies, for $k<n$:
$0=\int_{-1}^{1}T_{n}(t)T_{k}(t)\frac{1}{\sqrt{1-t^{2}}}dt=\frac{\pi}{n}\sum_{j=1}^{n+1}\lambda_{j}T_{n}(\xi_{j})T_{k}(\xi_{j}),$
@@
$=$ $\frac{1}{n}\sum_{j=1}^{n+1}\lambda_{j}(-1)^{1+j}\frac{1}{2}\left(T_{|k_{1}-k_{2}|}(\xi_{j})+T_{k_{1}+k_{2}}(\xi_{j})\right)$
@@ 5. Detecting the diagonally-determined property
The quadratic module of $g_{1},\ldots,g_{m}$, denoted by $\mathcal{Q}(g_{1},\ldots,g_{m})$, is defined as the set of all polynomials of the form
@@
We assume that $\mathcal{Q}(g_{1},\ldots,g_{m})$ satisfies the Archimedean condition, meaning that there exists $p\in\mathcal{Q}(g_{1},\ldots,g_{m})$ such that
$\left\{{\mathbf{x}}\in\mathbb{R}^{d}\;:\;p({\mathbf{x}})\geq 0\right\}\mbox{ is compact.}$
@@
[…]
@@
We assume that we now know $a,b]=\mathrm{diag}(\Omega)$, and that $a<0<b$. The latter assumption is without loss of generality, due to Lemma 2.9. Then, to detect whether the second property holds in Definition 2.1, consider the optimization problem:
@@
$\rho=\min_{{\mathbf{v}}^{\prime}}\,\{\,1-\langle{\mathbf{v}}^{\prime},\mathbf{1}\rangle:\>\ell_{{\mathbf{v}}^{\prime},a}({\mathbf{x}}),\ell_{{\mathbf{v}}^{\prime},b}({\mathbf{x}})\geq 0\quad\,\forall{\mathbf{x}}\in\Omega\,\}\,.$
@@
Assume $\Omega$ is of the form (5.1), that Assumption 5.1 holds, and that $\mathrm{diag}(\Omega)=[a,b]$ where $a<0<b$. Then one has $\rho_{k}\in[0,1]$ for all $k$. Moreover, if $\Omega$ is diagonally-determined, then the non-increasing sequence $\{\rho_{k}\}$ converges to zero. If $\rho_{k}=0$ for some finite value of $k$, then the optimal ${\mathbf{v}}^{\prime}\in\mathbb{R}^{d}$ of (5.4) certifies that $\Omega$ is diagonally-determined with vector ${\mathbf{v}}={\mathbf{v}}^{\prime}$ and diagonal $[a,b]$.
@@
where we again use that $a<0<b$. Thus, ${\mathbf{x}}\mapsto\langle(1-\varepsilon){\mathbf{v}},{\mathbf{x}}\rangle-a\in\mathcal{Q}(g_{1},\ldots,g_{m})$ and ${\mathbf{x}}\mapsto b-\langle(1-\varepsilon){\mathbf{v}},{\mathbf{x}}\rangle\in\mathcal{Q}(g_{1},\ldots,g_{m})$, by Theorem 5.2. This implies that, for sufficiently large $k^{*}\in\mathbb{N}_{0}$, ${\mathbf{v}}^{\prime}=(1-\varepsilon){\mathbf{v}}$ is feasible for (5.4) with $k=k^{*}$. Thus $\rho_{k^{*}}\leq 1-\langle(1-\varepsilon){\mathbf{v}},\mathbf{1}\rangle=\varepsilon$, and therefore the sequence $\{\rho_{k}\}$ is non-increasing, and converges to zero. ∎
@@
By Proposition 2.5, and its proof, we know that $\Omega$ is diagonally-determined, with diagonal [a,b]=[1,1][a,b]=[-1,1] and vector ${\mathbf{v}}=\frac{1}{d}\mathbf{1}$. We may describe $\Omega$ as a semi-algebraic set in the form (5.1) by using the polynomials $g_{j}({\mathbf{x}})=1-x_{j}^{2}$ for $j=1,\ldots,d$, i.e. $m=d$. The next step is to find the diagonal of $\Omega$ by finding the roots of the univariate polynomials $p_{j}(t):=g_{j}(t\mathbf{1})=1-t^{2}$ for $j=1,\ldots,d$. These polynomials are the same for all $j$ and nonnegative on the interval [1,1][-1,1], so that we obtain [a,b]=[1,1][a,b]=[-1,1]. The next step is to solve (5.4) for increasing values of $k\in\mathbb{N}$. Using the identity
$1\pm x=\frac{1}{2}\left((1\pm x)^{2}+1-x^{2}\right),\quad x\in\mathbb{R},$
@@
$\langle{\mathbf{v}},{\mathbf{x}}\rangle-a=\frac{1}{2d}\sum_{j=1}^{d}\left((1+x_{j})^{2}+g_{j}({\mathbf{x}})\right)\in\mathcal{Q}(g_{1},\ldots,g_{d})_{2}.$
@@ 6. Conclusion and discussion
The extremal problem (1.3) we considered in this paper relies on a specific generalization of the concept of a monic univariate polynomial to the multivariate case. In particular, in the multivariate case we require that the leading coefficients (in the standard monomial basis) sum to one, leading to the introduction of the set $\Pi^{*}_{n}$ in Definition 1.1. There are other natural choices, like also requiring the leading coefficients to be nonnegative. Having said that, the least Chebyshev polynomial in Theorem 3.1 already has nonnegative leading coefficients for many diagonally-determined sets, e.g. the unit ball in a monotone norm $\|\cdot\|$. To see this, let $\Omega$ be such a unit ball, centered at the origin. By Proposition 2.5, $\Omega$ is then diagonally-determined with diagonal $a,b]=[-1/\|\mathbf{1}\|,1/\|\mathbf{1}\|]$ and allows an entrywise-nonnegative vector $\mathbf{v}$}. By Theorem 3.1, the least Chebyshev polynomial in this case is given by
@@
A second point to mention, is that we have not explored the geometric and invariance properties of diagonally-determined sets to any great extent in this paper, and this deserves to be investigated further. The same may be said for the decision problem of whether a given set is diagonally-determined. If $\Omega$ is a compact semi-algebraic set, one may formulate an algorithmic procedure to decide this question, as we did in Section 5. An interesting special case is the (complexity) question: may one decide in polynomial time if a given polyhedron is diagonally-determined?